OpenStudy (anonymous):

How do you solve: Y=-3 Y=-2x+1 And put it on a graph, finding the solution?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

-7

OpenStudy (the_fizicx99):

y = -3 >.< plot a straight line on the y axis at (0,-3) then plot the y intercept of y = -2x + 1 -- (0,1) and apply rise over run. Where they intersect is your solution .-.

geerky42 (geerky42):

Since we alread knew that y = -3, plug that to second equation: y = -2x + 1 -3 = -2x + 1 solve for x then that it.

OpenStudy (the_fizicx99):

http://prntscr.com/3aeewg (2,-3) is your solution, when you plug in -3 for y into y = -3 you get \(\ -3 = -3 \color{green}{\checkmark} \) and \(\ - 3 = -2(2) + 1 \longrightarrow -3 = -4 + 1 \longrightarrow -3 = -3 \color{green}{\checkmark} \)